# Monte Carlo Tree Diffusion for System 2 Planning

### v2

@@ A.2 Heuristic Controller, Value-Learning Policy and Inverse Dynamics Model
A long‐standing challenge in diffusion‐based planning is the difficulty of preserving global trajectory coherence while managing local, high‐dimensional state‐action control (Chen et al., 2024a, b). For example, PlanDQ (Chen et al., 2024b) couples a high‐level diffusion planner with a learned low‐level policy. Similarly, our approach focuses the diffusion planner on lower‐dimensional, representative state information (e.g., object or agent positions), while delegating detailed action inference to the hueristic controller or value-learning policy or inverse dynamics model. The details of this integration is discussed in Algorithm 10. Such a hierarchical design allows MCTD and baseline planners to concentrate on broader strategic decisions while leaving the finer‐grained execution details to the heuristic controller or value-learning policy or inverse dynamics model.
@@ A.3.1 Diffuser
Random Search Variant: To evaluate the impact of test-time compute scaling with more samples, we implemented Diffuser version ”Sample-Score-Rank” (Zhou et al., 2024). Additionally, each sample can be guided through different guidances, we applied this by uniformly distributing the guidance set to the samples.
@@ A.3.4 Value-Learning Policy
For some tasks, to handle complex action space, we applied a value-learning policy by modifying the PlanDQ (Chen et al., 2024b) source code, https://github.com/changchencc/plandq.
@@ A.4 The Evaluation Details
We followed the OGBench (Park et al., 2024) task configurations, 5 tasks per each environment. For each task, we evaluated 10 random seeds per each model, and reported the averaged success rates and their standard deviations.
@@ A.5.1 Modifications from OGBench
Although the original OGBench datasets include random noise in start and goal positions, we remove this noise to isolate model performance from environmental randomness. This modification facilitates clearer comparisons across methods and ensure that differences in performance can be attributed to planning capabilities rather than uncontrolled stochasticity. Additionally, we incorporate velocity into the state representation, which the original benchmark does not provide, in order to apply the heuristic controller for pointmaze tasks implemented in (Janner et al., 2022). For antmaze giant task, we found there are some cases where the plans are well generated, but due to the inaccurate execution, it cannot be succeed. To prevent this, we extend the episode length from to .
@@ A.5.2 Guidance Function
To plan long-horizon while sustaining the plan effectiveness to reach the goal as fast as possible, we applied the guidance function to reduce the distances between every state and goal, (Chen et al., 2024a). We note that, to do fair comparison we applied the same guidance style over every baseline. For example, Diffuser originally used the goal-inpainting guidance for pointmaze, but in this evaluation, we used the distance guidance.
@@ A.5.5 Heuristic Controller and Value-Learning Policy
For pointmaze, we take the heuristic controller designed in (Janner et al., 2022). For antmaze, we use the value-learning policy (Wang et al., 2022) as done in (Chen et al., 2024b). However, we note that different from (Chen et al., 2024b), we only give the steps further planned position information as the subgoal for that.
@@ A.5.8 Comparison with Goal‐Inpainting Diffuser
Diffuser (Janner et al., 2022) typically employs a learned guidance function to bias generated trajectories toward high‐return outcomes. For pointmaze tasks, however, an alternative approach—goal‐inpainting guidance—has also been introduced (Janner et al., 2022), which further improves single‐pass planning by “imagining” an intermediate trajectory whose final state is the designated goal. In the context of Diffuser‐Replanning, we adapt goal‐inpainting by providing the experienced trajectory as contextual information rather than a direct target for denoising.
@@ A.6.6 Value-Learning Policy
Similar to antmaze tasks, we applied value-learning policy (Wang et al., 2022) by achieving the subgoals suggested by planners for steps further planned states.
@@ A.7.1 Setup
Because diffusion‐based planners typically operate on state‐action representations, we first learn an image encoder via a Variational Autoencoder (VAE) (Kingma & Welling, 2013). The resulting latent states serve as inputs to the diffusion models. Notably, these latent encodings do not include explicit positional information, so the planner must infer underlying geometry and navigate the maze indirectly. We design an inverse dynamics model (a small network of linear layers) to convert consecutive latent observations into actions: . This replaces the heuristic controller of the original Diffuser, which cannot be directly applied here due to partial observability.
@@ A.7.4 Dataset Generation and Planning Horizon
We follow the OGBench (Park et al., 2024) dataset generation script for PointMaze, with trajectory horizons of 1000 for Medium and Large mazes. To enhance training, we use reduced planning horizons (500 for Medium/Large) and apply a sliding-window technique, as in our earlier pointmaze setup.
@@ A.7.5 Variational Autoencoder
We pretrained a VAE (Kingma & Welling, 2013) to encode RGB observations into an -dimensional latent space, compressing high-dimensional inputs into a structured representation for downstream planning. It represents the latent variable as a Gaussian distribution with mean and standard deviation , using the reparameterization trick:
@@ Appendix B MCTD Architecture Implementation
Next, we adopt a causal noise schedule (Chen et al., 2024a), which renders the denoising process semi‐autoregressive: earlier plan segments undergo more thorough denoising, while subsequent ones shift more gradually from higher to lower noise. This preserves the globally consistent trajectory generation characteristic of diffusion models but also ensures that critical decisions near the beginning of the plan receive higher‐fidelity refinements. Finally, both planning and simulation employ partial and jumpy denoising via DDIM sampling (Song et al., 2020a), allowing the system to skip directly from noise level to . This skip‐level update accelerates the evaluation of candidate trajectories during tree expansion and supports partial denoising steps essential for exploration.

### v3

@@ A.2 Heuristic Controller, Value-Learning Policy and Inverse Dynamics Model
A long‐standing challenge in diffusion‐based planning is the difficulty of preserving global trajectory coherence while managing local, high‐dimensional state‐action control (diffusion_forcing; plandq). For example, PlanDQ (plandq) couples a high‐level diffusion planner with a learned low‐level policy. Similarly, our approach focuses the diffusion planner on lower‐dimensional, representative state information (e.g., object or agent positions), while delegating detailed action inference to the hueristic controller or value-learning policy or inverse dynamics model. The details of this integration is discussed in Algorithm 10. Such a hierarchical design allows MCTD and baseline planners to concentrate on broader strategic decisions while leaving the finer‐grained execution details to the heuristic controller or value-learning policy or inverse dynamics model.
@@ A.3.1 Diffuser
Random Search Variant: To evaluate the impact of test-time compute scaling with more samples, we implemented Diffuser version ”Sample-Score-Rank” (d-mpc). Additionally, each sample can be guided through different guidances, we applied this by uniformly distributing the guidance set to the samples.
@@ A.3.4 Value-Learning Policy
For some tasks, to handle complex action space, we applied a value-learning policy by modifying the PlanDQ (plandq) source code, https://github.com/changchencc/plandq.
@@ A.4 The Evaluation Details
We followed the OGBench (4_park2024ogbench) task configurations, 5 tasks per each environment. For each task, we evaluated 10 random seeds per each model, and reported the averaged success rates and their standard deviations.
@@ A.5.1 Modifications from OGBench
Although the original OGBench datasets include random noise in start and goal positions, we remove this noise to isolate model performance from environmental randomness. This modification facilitates clearer comparisons across methods and ensure that differences in performance can be attributed to planning capabilities rather than uncontrolled stochasticity. Additionally, we incorporate velocity into the state representation, which the original benchmark does not provide, in order to apply the heuristic controller for pointmaze tasks implemented in (diffuser). For antmaze giant task, we found there are some cases where the plans are well generated, but due to the inaccurate execution, it cannot be succeed. To prevent this, we extend the episode length from to .
@@ A.5.2 Guidance Function
To plan long-horizon while sustaining the plan effectiveness to reach the goal as fast as possible, we applied the guidance function to reduce the distances between every state and goal, (diffusion_forcing). We note that, to do fair comparison we applied the same guidance style over every baseline. For example, Diffuser originally used the goal-inpainting guidance for pointmaze, but in this evaluation, we used the distance guidance.
@@ A.5.5 Heuristic Controller and Value-Learning Policy
For pointmaze, we take the heuristic controller designed in (diffuser). For antmaze, we use the value-learning policy (dql) as done in (plandq). However, we note that different from (plandq), we only give the steps further planned position information as the subgoal for that.
@@ A.5.8 Comparison with Goal‐Inpainting Diffuser
Diffuser (diffuser) typically employs a learned guidance function to bias generated trajectories toward high‐return outcomes. For pointmaze tasks, however, an alternative approach—goal‐inpainting guidance—has also been introduced (diffuser), which further improves single‐pass planning by “imagining” an intermediate trajectory whose final state is the designated goal. In the context of Diffuser‐Replanning, we adapt goal‐inpainting by providing the experienced trajectory as contextual information rather than a direct target for denoising.
@@ A.6.6 Value-Learning Policy
Similar to antmaze tasks, we applied value-learning policy (dql) by achieving the subgoals suggested by planners for steps further planned states.
@@ A.7.1 Setup
Because diffusion‐based planners typically operate on state‐action representations, we first learn an image encoder via a Variational Autoencoder (VAE) (vae). The resulting latent states serve as inputs to the diffusion models. Notably, these latent encodings do not include explicit positional information, so the planner must infer underlying geometry and navigate the maze indirectly. We design an inverse dynamics model (a small network of linear layers) to convert consecutive latent observations into actions: . This replaces the heuristic controller of the original Diffuser, which cannot be directly applied here due to partial observability.
@@ A.7.4 Dataset Generation and Planning Horizon
We follow the OGBench (4_park2024ogbench) dataset generation script for PointMaze, with trajectory horizons of 1000 for Medium and Large mazes. To enhance training, we use reduced planning horizons (500 for Medium/Large) and apply a sliding-window technique, as in our earlier pointmaze setup.
@@ A.7.5 Variational Autoencoder
We pretrained a VAE (vae) to encode RGB observations into an -dimensional latent space, compressing high-dimensional inputs into a structured representation for downstream planning. It represents the latent variable as a Gaussian distribution with mean and standard deviation , using the reparameterization trick:
@@ Appendix B MCTD Architecture Implementation
Next, we adopt a causal noise schedule (diffusion_forcing), which renders the denoising process semi‐autoregressive: earlier plan segments undergo more thorough denoising, while subsequent ones shift more gradually from higher to lower noise. This preserves the globally consistent trajectory generation characteristic of diffusion models but also ensures that critical decisions near the beginning of the plan receive higher‐fidelity refinements. Finally, both planning and simulation employ partial and jumpy denoising via DDIM sampling (ddim), allowing the system to skip directly from noise level to . This skip‐level update accelerates the evaluation of candidate trajectories during tree expansion and supports partial denoising steps essential for exploration.